When dealing with imaginary numbers, how do you simplify?
I^12
I know the answer is 1 but how do I get the answer with remembering the cycle?

Answer by Alan3354(69443)

i^12
--------
i^2 = -1
i^3 = -i
i^4 = 1
--------
i^12 = (i^4)^3 = 1^3
= 1